OpenStudy (anonymous):

MEDAL AND FAN A system of equations is shown below: 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

@lizz123

OpenStudy (anonymous):

@alauren

OpenStudy (anonymous):

@deh_vinn

OpenStudy (anonymous):

@DemiUnicorn

OpenStudy (anonymous):

@amoodarya

OpenStudy (anonymous):

@DanJS

OpenStudy (amoodarya):

8x + 5y = 9 a 3x + 2y = 4 b so a+b= 8x+5y+3x+2y=9+4 11x+7y=13 now 11x=7y=13 3x+2y=4 it is a new system solve it

OpenStudy (anonymous):

sorry could you help me out im really bad with math